OpenStudy (anonymous):

express as a single logarithm: log(x/y^2)+3logy-2logxy^3

OpenStudy (anonymous):

\[\log \frac{ x }{ y ^{2} }+3\log y-2\log xy ^{3}\]

OpenStudy (callisto):

\[\log\frac{x}{y^2}+3\log y-2\log xy^3\]\[=\log\frac{x}{y^2}+\log y^3-\log (xy^3)^2\]\[=\log(\frac{x}{y^2}\times y^3 \div (xy^3)^2) =...\]

OpenStudy (callisto):

\[a \log x = \log x^a\]\[\log m + \log n =\log (mn)\]\[\log m -\log n = \log (\frac{m}{n})\]

OpenStudy (anonymous):

thanks. can you answer one last question?

OpenStudy (anonymous):

after this one

OpenStudy (anonymous):

sorry bad joke

OpenStudy (anonymous):

\[\log_{6}(x+5)=\frac{ 1 }{ 2 } \] I think it's not possible because x+5 would have to be negative but not sure